OpenStudy (anonymous):

Evaluate 5 - 3/2 to the 3rd power

OpenStudy (compassionate):

\[(5 - \frac{ 3 }{ 2 })^3\] Multiply everything to the third power

OpenStudy (photon336):

@IsaiahMoore14 you need to subtract first 5-(3/2)

OpenStudy (photon336):

PEMDAS. we do what's in the ( ) first and that's subtraction (5-3/2)^(3) then afterwards, we raise (3.5)^3 that just means \[5-\frac{ 3 }{ 2 } = 3.5 \] \[(3.5)^{3} = (3.5)(3.5)(3.5) = 42.875\]